Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Independent Events

Expert replies
by manik11 » Thu Nov 19, 2015 4:18 am
Events A and B are independent and have equal probabilities of occurring. What is the probability that event A occurs?

1) The probability that at least one of events A and B occurs is .84.
2) The probability that event B occurs and event A does not is .24.

OA : A
Source : Veritas Prep

Even after going through the official explanation I can't understand why statement 2 is insufficient.
I also often get confused when the phrase "at least" is used in the probability questions.
Experts...could you guys please help me with this one?
Join the discussion
Source: — Data Sufficiency |

by GMATGuruNY » Thu Nov 19, 2015 4:51 am
manik11 wrote:Events A and B are independent and have equal probabilities of occurring. What is the probability that event A occurs?

1) The probability that at least one of events A and B occurs is .84.
2) The probability that event B occurs and event A does not is .24.
A and B have equal probabilities of occurring.
Thus:
A = B.
not A = not B.

P(at least A or B) + P(neither A nor B) = 1.
Put another way:
P(neither A nor B) = 1 - P(at least A or B).

Statement 1: The probability that at least one of events A and B occurs is .84.
Thus:
P(neither A nor B occurs) = 1 - 0.84 = 0.16.
(not A)(not B) = 0.16.

Since not B = not A, we get:
(not A)(not A) = 0.16
not A = 0.4.

Thus, A = 1 - 0.4 = 0.6.
SUFFICIENT.

Statement 2: The probability that event B occurs and event A does not is .24.
In other words:
(B)(not A) = 0.24.
Since B = A, we get:
(A)(not A) = 0.24.

This equation is satisfied by the values yielded in Statement 1.
If A = 0.6 and not A = 0.4, then (A)(not A) = (0.6)(0.4) = 0.24.
But this equation is also satisfied if A = 0.4 and not A = 0.6:
(A)(not A) = (0.4)(0.6) = 0.24.

Since A = 0.6 in the first case but A = 0.4 in the second case, INSUFFICIENT.

The correct answer is A.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by [email protected] » Thu Nov 19, 2015 9:50 am
Hi manik11,

We're told that events A and B are independent and have EQUAL PROBABILITY of occurring. We're asked for the probability that event A occurs.

Fact 1: The probability that AT LEAST one of the events occurs is .84

This means that A, or B or BOTH occurs = .84

We can restate this to mean that the probability that neither A, nor B nor BOTH occurs = .16

(Not A)(Not B) = .16

Since we're told that A and B have equal probabilities of occurring, they also have equal probabilities of NOT occurring. So….

(Not A) = (Not B)

Combining these 2 equations gives us the probability that A will not occur…..

(.4)(.4) = .16

The probability that A DOES NOT occur = .4
The probability that A DOES occur = .6
Fact 1 is SUFFICIENT

Fact 2: The probability that B occurs and A DOES NOT occur = .24

This tells us that….

(B)(Not A) = .24

From the prompt, we know that…

A = B
(Not A) = (Not B)
A + (Not A) = 1

In this scenario, the values are .4 and .6, but we don't know which is which (one is the probability the event occurs, the other is the probability that the event does NOT occur). Either one COULD be the answer.
Fact 2 is INSUFFICIENT.

Final Answer: A

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion

by Max@Math Revolution » Sun Nov 22, 2015 1:10 am
Forget conventional ways of solving math questions. In DS, Variable approach is the easiest and quickest way to find the answer without actually solving the problem. Remember equal number of variables and independent equations ensures a solution.

Events A and B are independent and have equal probabilities of occurring. What is the probability that event A occurs?

1) The probability that at least one of events A and B occurs is .84.
2) The probability that event B occurs and event A does not is .24.

There are 3 variables ( P(A), P(B), P(A&B)), and 2 equations (P(A)=P(B), P(A and B)=P(A)(B)), and 2 more equations are given from the 2 conditions, so there is high chance (D) will be the answer.
For condition 1, P(A or B)=P(A)+P(B)+P(A and B)=2P(A)-P(A)^2=0.84, P(A)^2-2P(A)+0.84=0. (P(A)-0.6)(P(A)-1.4)=0, and P(A)=0.6, 1.4, but probability cannot be greater than 1, so P(A)=0.6, and this is sufficient.
For condition 2, inP( only B)=P(B)-P(A and B), but we cannot know the value of P(B), so this is insufficient, and the answer becomes (A).

For cases where we need 1 more equation, such as original conditions with “1 variable”, or “2 variables and 1 equation”, or “3 variables and 2 equations”, we have 1 equation each in both 1) and 2). Therefore, there is 59 % chance that D is the answer, while A or B has 38% chance and C or E has 3% chance. Since D is most likely to be the answer using 1) and 2) separately according to DS definition. Obviously there may be cases where the answer is A, B, C or E.
Join the discussion